OpenStudy (anonymous):

Which equation is in point-slope form and is depicted by the line in this graph? a.(y-3)=3/5(x-2) b.(y-3)=-3/5(x-2) c.(y-3)=5/3(x-2) d. (y+3)=5/3(x+2)

OpenStudy (anonymous):



OpenStudy (anonymous):

@Janu16

OpenStudy (janu16):

just write all of them in point slope and try graphing them

OpenStudy (anonymous):

I don't have a graphing calculator

OpenStudy (anonymous):

@texaschic101

OpenStudy (janu16):

use desmos.com

OpenStudy (texaschic101):

slope = (0 - 3) / (3 + 2) = - 3/5 y - y1 = m (x - x1)....(- 2,3) y - 3 = - 3/5 (x + 2)